Errata to Mathematics for the Analysis of Algorithms, third edition (last updated 07 July 2018)

Page 1, eq. (1.1) integer n integer n ≥ 0 Page 16, in (2.27)

∑_{k=1}^{n−3} ∑_{k=0}^{n−3}

Page 25, after the theorem [Further results on such recurrences have been obtained by T. R. Walsh, Information Processing Letters 19 (1984), 203–208; S. Kapoor and E. M. Reingold, J. Math. Analysis and Applications 109 (1985), 591–604.] Page 25, in (2.64) A_{h+1} A_{h+1}(z) Page 26, line 6 we were lead we were led Page 48, line 2 from the bottom in equation (4.22) in (4.22) Page 77–80, throughout the bibliography adding several commas would make the format more consistent Page 79, line 25 leurs relation leurs relations Page 80, bottom line 5(1) 5(3) Page 87, line 7 (1 1 Page 90, line 8 from the bottom 0676 0739 Page 94, add a new (one-line) paragraph Curiously, (D19) is wrong! Page 104, line 14 from the bottom z_{1} … z_{n−k} z_{2} … z_{n−k} Page 106, bottom of page [Another solution to this problem, which leads to a generating function, has been published by P. V. Poblete, Journal of Algorithms 4 (1983), 388–393.] Page 132, line 26 of the right-hand column Louis Xavier Joseph Jean Gustave Nicolas